OpenStudy (anonymous):

what is the slope of the line that passes thru the points (-3, -10) & (5,-7) ?

OpenStudy (anonymous):

first y value minus the secound divided by the 1st x value minus the second

OpenStudy (anonymous):

Slope ,m=y-y1/(x-x1) Here y=-10 y1=-7 x=-3 x1=5

OpenStudy (anonymous):

There fore m=-10-(-7)/(-3-5)

OpenStudy (anonymous):

=-3/-8=3/8